A demand function is given as Qx=500-2Px+10Y+3Py where Qx=Quantity demanded for x goods
Px=Price of X goods
Py=Price of Y goods
Y=Income of consumer
Find price elasticity of demand at price 50
Find income elasticity of demand at price 1000
Find cross elasticity of demand of Y goods at price 40
1
Expert's answer
2019-06-10T09:32:27-0400

Price elasticity of demand at price 50 is:

Ed = - 2×50/Qx.

Income elasticity of demand at price 1000 is:

Ed = 10×1000/Qd.

Cross elasticity of demand of Y goods at price 40 is:

Ed = 3×40/Qd.